At BCD car wash, when a customer arrives, he or she first checks in with a receptionist and then leaves the car to one of several car washers. While the customer is waiting for the car to be washed, he or she proceeds to pay at the cash register. After the payment, the customer returns to his/her car when it is done and then leaves.

On average, reception takes 3 minutes. Car wash takes 5 minutes. Paying at the cash register takes 4 minutes. Returning to car and leaving requires 2 minutes. Currently, the car wash employs 1 receptionist, 2 cash registers and 3 car washers.

What is the bottleneck of the process?

I.

Receptionist

II.

Cash register

III.

Car washer

IV.

Both cash register and car washer
Answer #1

Capacity per hour = Number of employees*60/Service time

Reception = 1*60/3 = 20

Cash register = 2*60/4 = 30

Car wash = 3*60/5 = 36

Returning and leaving = 1*60/2 = 30

Activities of cash register and car wash is in parallel but The bottleneck process is receptionist due to least capacity

Ans: 1 (Receptionist)